OpenStudy (anonymous):

Calculus: Need help on a tangent line question Find the equation of the tangent line to the curve at the given point x=cos(7y); ((-1/2)sqrt(3), (-1/6)pi)

OpenStudy (wolfe8):

So the curve is ((-1/2)sqrt(3), (-1/6)pi) ? Do you know what the derivative of a curve equation gives you?

OpenStudy (anonymous):

I got the derivative but im not sure if it is correct but \[1=-7\sin7y \frac{ dy }{ dx }\]

OpenStudy (anonymous):

I keep trying to plug to solve for dy dx but I end up with 1/-24.5

OpenStudy (anonymous):

But the awnser choices are

OpenStudy (wolfe8):

Ok first of all I got that wrong. What I stated was the point. So for the equation, you'll have to solve for y and then derive. Let me see.

OpenStudy (wolfe8):

I get d/dx(1/7 cos^(-1)(x))

OpenStudy (anonymous):

How did you get to that step?

OpenStudy (wolfe8):

DANGIT MY TYPING TODAY. What I put was the rearranged equation. Here lemme show you my steps from the beginning. First I rearranged it in terms of y and get \[y=\frac{ \cos^{-1} x }{ 7 }\] Then I differentiate it. I afctor out the constant 1/7 and use the derivative of cos^(-1)(x) is -1/sqrt(1-x^2).

OpenStudy (wolfe8):

So the final derivative is \[\frac{ 1 }{ 7 }\frac{ -1 }{ \sqrt{1-x ^{2}} }\]

OpenStudy (anonymous):

Ok so now I just plug in x and I get \[(\frac{ 1 }{ 7 })(\frac{ -1 }{\sqrt{1-\frac{ \sqrt{3} }{ 2 }^2}}\]

OpenStudy (wolfe8):

I'm sure you can simplify that further. The square root of 3 is squared :)

OpenStudy (anonymous):

is just 3/4

OpenStudy (anonymous):

So the slope is (-4/7)

OpenStudy (wolfe8):

Hold up. Why does the picture you posted not look like the one we're doing here

OpenStudy (anonymous):

Because it is the missed problem I got and I needed help on why I got it wrong

OpenStudy (anonymous):

or am doing wrong

OpenStudy (wolfe8):

Well which one are you asking for help for right now? I've been doing the original one you posted. Also, you put in x wrong. Read again what given x you have. I think you just put in sqrt(3) instead of the whole thing.

OpenStudy (anonymous):

They both are the same

OpenStudy (anonymous):

They both are the same problem

OpenStudy (wolfe8):

Hmm. Well for the original one check again your work when pluggin in the x as I said. That will give you the slope. Now, for the second part, do you remember the slope of a horizontal line?

OpenStudy (anonymous):

Yes you just set f(x)=0

OpenStudy (anonymous):

correct?

OpenStudy (anonymous):

or wait woops wrong problem on the picture

OpenStudy (wolfe8):

Not quite. You set f'(x)=0 It is the derivative, the slope that is 0

OpenStudy (anonymous):

Lol sorry I dont need help on that one my bad

OpenStudy (anonymous):

I uploaded the wrong one

OpenStudy (anonymous):

Now im tripping lol

OpenStudy (wolfe8):

God my head hurts too now. Well let's continue where we left off then. Plug in x, correctly.

OpenStudy (anonymous):

ok

OpenStudy (dan815):